40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 67     Entry time: Tue Nov 6 10:42:01 2007
Author: rob 
Type: Configuration 
Category: IOO 
Subject: mode cleaner locked 
Increased the power exiting the PSL by turning the half-wave plate after the MOPA, opened the PSL shutter, and aligned the mode cleaner to the input beam. It wasn't that hard to find the beam with the aperture open all the way on the MC2 camera. The transmitted power is now 2.9 arbitrary units, while the input power is 1.2 arbitrary units. Not sure yet if that's an increase or decrease in efficiency, since no one posted numbers before the vent. Also turned on the input-steering PZTs and saw a REFL beam on the camera.
ELOG V3.1.3-